# PEFT-Bench: A Parameter-Efficient Fine-Tuning Methods Benchmark

### v1

@@ 5 PEFT-Bench Results
Figure 4: Bar chart showing the stability of different PEFT methods on 4 low-resource datasets. IA3 achieves the lowest standard deviation on average across all datasets, and LoRA is less stable than other methods with CB datasets, while also achieving a better average score. Additionally, P-Tuning generally achieves the worst results in terms of performance and stability. For numerical results please see Table 11 in Appendix D.
LoRA achieves better performance, but LNTuning is more efficient.
From the performance evaluation results in Tables 1 and 2, we can see that LoRA achieves the highest performance on many tasks. However, we can also see that a more parameter-efficient method, LNTuning, that trains a smaller number of parameters, was able to achieve better performance in MNLI, SST-2, and MMLU datasets. This efficiency difference was captured by our PSCP metric in Table 3 as it balanced the final result and offset the LNTuning method for the lower number of parameters (compared to LoRA).
@@
The results in Tables 1 and 2 show that the soft prompt-based methods perform generally worse compared to other methods, and in some cases, they achieve an F1 score of 0, mostly because they collapse into predicting only a single class. To achieve better results, more fine-grained per-dataset hyperparameter fine-tuning would probably be required. As the out-of-the-box usability and versatility across different datasets are also important factors when designing an efficient method.
In addition, we have been witnessing memory spikes when training the Prompt Tuning method. These spikes usually happened during the first 1000 steps of training, and they were much less noticeable when training on small amounts of data. After that, the memory usage was stable for the rest of the training. We believe this may be caused by the direct addition of a soft prompt into the input of an autoregressive model.
@@
Finally, we also evaluate the stability of different PEFT methods. Since stability is not directly associated with the efficiency of PEFT methods, we decided to create a separate category for our evaluation. The results of the stability experiments can be seen in Figure 4. We evaluated the stability of individual PEFT methods on both, low (CB, COPA, SVAMP, CoLA, and WSC) as well as high resource (SST-2) datasets. Based on the results, the method that achieved the lowest standard deviations on average was IA3. Moreover, we can see that most of the evaluated PEFT methods achieve relatively stable results on small data, except for LoRA on the CB and WSC dataset, where the standard deviations are larger compared to other methods.
@@ 6 Conclusion
In our work, we introduce PEFT-Bench, the first unified and comprehensive benchmark for parameter-efficient fine-tuning methods for NLP, and PEFT-Factory, as an underlying technological framework supporting PEFT-Bench. PEFT-Bench includes 27 datasets from various NLP problems and domains, implements support for off-the-shelf PEFT methods, simplifies the evaluation of new custom PEFT methods, and finally, introduces 6 PEFT methods as ready-to-use state-of-the-art baselines. We also propose the PSCP metric to properly incorporate efficiency factors like parameters, inference time, and memory usage into the evaluation.
In our comparison, we show that performance on downstream tasks should not be the only condition when evaluating PEFT methods, and sometimes a more efficient PEFT method, like LNTuning, can compensate for the performance with its efficiency. Additionally, we experienced that the soft prompt-based methods are generally harder to train and require extensive hyperparameter search.
We view PEFT-Bench as a long-term evolving project, which will be continuously updated with new features and the addition of new evaluations for different PEFT methods. In the near future, we plan to deploy a web interface that would allow viewing of the results and analysis of PEFT-Bench. Additionally, we also plan to use this benchmark to evaluate multi-task learning factors that affect the performance of multi-task PEFT methods.
@@ Acknowledgments
This work was partially funded by the European Union NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the projects No. 09I01-03-V04-00006 and 09I01-03-V04-00068; by the European Union, under the project LorAI - Low Resource Artificial Intelligence, GA No. 101136646.
@@
We also run PEFT-Bench on a single model – LLaMa3-8B-Instruct. This may potentially limit the generalizability of PEFT-Bench, but similar to language, PEFT methods are not really dependent on the model architecture, and their performance would generally correlate with the performance of the model.
Finally, we evaluate 6 different PEFT methods, with the majority being from the additive PEFT and soft prompt-based category. To keep experiments computationally feasible, we rely on the evaluated PEFT methods with limited hyperparameter tuning. Thanks to PEFT-Factory, the whole benchmark is easily extendable and can add their own PEFT method(s).
@@
The experiments in this paper were conducted with publicly available datasets MNLI, QQP, QNLI, SST2, STS-B, MRPC, RTE, CoLA, MultiRC, ReCoRD, BoolQ, WiC, WSC, CB, COPA, MMLU, PIQA, SIQA, HellaSwag, WinoGrande, OpenBookQA, MathQA, GSM8K, SVAMP, Conala, CodeAlpacaPy, and APPS, citing the original authors. As we were not able to determine the license for all used datasets, we have opted to use them in the limited form possible, adhering to the terms of use of the GLUE and SuperGLUE benchmarks. As the datasets are commonly used in other related works and were published in scientific works that went through an established review process, we do not check for the presence of any offensive content, as it was already removed by the authors of these publicly available datasets. In addition, we do not utilize any personally identifiable information or offensive content, and we do not perform crowdsourcing in any form for data annotation. To our knowledge, we are not aware of any potential ethical harms or negative societal impacts of our work, apart from the ones related to the field of Machine Learning (i.e., the use of computational resources that are consuming energy and producing heat with indirect CO2 emission production). We follow the license terms for the LLaMa-3-8B-Instruct model we use – all models and datasets allow their use as part of the research. As we transform conditional generation into the classification problem (generating only labels), in most cases, we minimize the problem of generating offensive or biased content.
@@
The experiments in this paper require a significant amount of GPU computing resources as we train and evaluate 1 model over multiple random initializations (5) for different methods (6) and datasets (27). Overall, the experiments, including evaluations (which did not require training, but still used GPU resources for inference) and preliminary experiments (which are not reported in the scope of our work), were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCO2eq/kWh. Approximately 3000 hours of computation were performed on hardware of type A100 PCIe 40GB (TDP of 250W). Total emissions are estimated to be 550.24 kgCO2eq, of which 0 percent were directly offset. Whenever possible, we tried to reduce the computational costs.
@@ Appendix B PSCP Importance Configurations
In this section, we provide an extended example of different configurations for importance $\beta$ for each efficiency factor. Based on the results from Table 9, we can see that with the change of importance for each parameter, the order of the results can change (e.g., in Table 9(a), the LoRA method achieves the best PSCP score, since we do not penalize the parameters cost as much as the other efficiency factors). Each sub-table represents a different use case, where we emphasize different efficiency factors. To also evaluate the changes in the magnitude of the importance $\beta$ we provide results with different magnitudes in Table 10. The order of the methods does not change, only the magnitude of penalized scores (unlike the results for different importance configurations in Table 9).
@@
Table 9: Results with PSCP metric under different importance settings. Notice how the order of the methods changes in Table (a) when we lower the $\beta_{p}$ importance factor.
PEFT Method $P_{avg}$ $cost_{p}^{-1}$ $cost_{f}^{-1}$ $cost_{m}^{-1}$ PSCP
IA3 74.7 1.0 0.97 0.77 55.62
Prompt Tuning 50.0 1.0 0.86 0.69 29.49
Prefix Tuning 45.9 0.94 1.0 0.77 33.18
P-Tuning 51.7 0.9 0.86 0.73 29.26
LoRA 80.1 0.97 1.0 0.77 60.08
LNTuning 77.8 1.0 1.0 0.77 60.19
PEFT Method $P_{avg}$ $cost_{p}^{-2}$ $cost_{f}^{-2}$ $cost_{m}^{-2}$ PSCP
IA3 74.7 1.0 0.94 0.59 41.41
Prompt Tuning 50.0 1.0 0.73 0.47 17.39
Prefix Tuning 45.9 0.88 1.0 0.6 23.98
P-Tuning 51.7 0.82 0.73 0.53 16.56
LoRA 80.1 0.94 1.0 0.6 45.06
LNTuning 77.8 1.0 1.0 0.6 46.57
PEFT Method $P_{avg}$ $cost_{p}^{-3}$ $cost_{f}^{-3}$ $cost_{m}^{-3}$ PSCP
IA3 74.7 1.0 0.91 0.45 30.83
Prompt Tuning 50.0 1.0 0.63 0.33 10.26
Prefix Tuning 45.9 0.82 1.0 0.46 17.33
P-Tuning 51.7 0.74 0.63 0.39 9.37
LoRA 80.1 0.92 1.0 0.46 33.8
LNTuning 77.8 1.0 1.0 0.46 36.03
PEFT Method $P_{avg}$ $cost_{p}^{-4}$ $cost_{f}^{-4}$ $cost_{m}^{-4}$ PSCP
IA3 74.7 1.0 0.89 0.35 22.95
Prompt Tuning 50.0 1.0 0.54 0.23 6.05
Prefix Tuning 45.9 0.77 1.0 0.36 12.53
P-Tuning 51.7 0.67 0.54 0.29 5.31
LoRA 80.1 0.89 1.0 0.36 25.35
LNTuning 77.8 1.0 1.0 0.36 27.88
Table 10: Results with PSCP metric under equal importance settings with different magnitudes. Notice how the order of the methods remains the same, but only the magnitude of the penalized score changes.

### v2

@@ 5 PEFT-Bench Results
Figure 4: Bar chart showing the stability of different PEFT methods on 4 low-resource datasets. IA3 achieves the lowest standard deviation on average across all datasets, and LoRA is less stable than other methods with CB datasets, while also achieving a better average score. Additionally, P-Tuning generally achieves the worst results in terms of performance and stability. For numerical results, see Table 11 in Appendix D.
LoRA achieves better performance, but BitFit and LNTuning are more efficient.
From the performance evaluation results in Tables 1 and 2, we can see that LoRA achieves the highest performance on many tasks. However, we can also see that a more parameter-efficient methods, BitFit and LNTuning, that train a smaller number of parameters and spend less GPU memory during training, were able to achieve better or on par performance in MNLI, SST-2, RTE and MMLU datasets. This efficiency difference was captured by our PSCP metric in Table 3, as it balanced the final result and offset the BitFit and LNTuning methods due to the lower number of parameters and lower memory consumption during training (compared to LoRA).
@@
The results in Tables 1 and 2 show that the soft prompt-based methods (i.e., Prompt Tuning, Prefix Tuning, and P-Tuning) perform generally worse compared to other methods, and in some cases, they achieve an F1 score of 0, mostly because they collapse into predicting only a single class. To achieve better results, more fine-grained per-dataset hyperparameter fine-tuning would probably be required. This suggests that out-of-the-box usability and versatility across different datasets are also important factors in designing an efficient method.
Additionally, we have observed memory spikes during training of the Prompt Tuning method. These spikes typically occurred during the first 1,000 steps of training, and they were significantly less noticeable when training on smaller amounts of data. After that, the memory usage was stable for the rest of the training. We believe this may be caused by the direct addition of a soft prompt into the input of an autoregressive model.
@@
Finally, we also evaluate the stability of different PEFT methods. Since stability is not directly associated with the efficiency of PEFT methods, we decided to evaluate it separately from the previously-described performance and efficiency results. The results of the stability experiments can be seen in Figure 4. We evaluated the stability of individual PEFT methods on both, low (CB, COPA, SVAMP, CoLA, SST-2, HellaSwag, and WSC) as well as high resource (SST-2) datasets. Based on the results, the method that achieved the lowest standard deviations on average was IA3. Moreover, we can see that most of the evaluated PEFT methods achieve relatively stable results on small data, except for LoRA on the CB and WSC dataset, where the standard deviations are larger compared to other methods.
@@ 6 Conclusion
In our work, we introduce PEFT-Bench, the first unified and comprehensive benchmark for parameter-efficient fine-tuning methods for NLP, and PEFT-Factory, as an underlying technological framework supporting PEFT-Bench. PEFT-Bench includes 27 datasets from various NLP problems and domains, implements support for off-the-shelf PEFT methods, simplifies the evaluation of new custom PEFT methods, and finally, introduces 7 PEFT methods as ready-to-use state-of-the-art baselines. We also propose the PSCP metric to properly incorporate efficiency factors like parameters, inference time, and memory usage into the evaluation.
In our comparison, we show that performance on downstream tasks should not be the only condition when evaluating PEFT methods, and sometimes, more efficient PEFT methods, like BitFit and LNTuning, can compensate for the performance with their efficiency. Additionally, we found that soft prompt-based methods are generally harder to train and require extensive hyperparameter tuning.
We view PEFT-Bench as a long-term, evolving initiative that will be continuously updated with new features and the addition of evaluations for different PEFT methods. The current state of PEFT-Bench, together with up-to-date results, is available at its GitHub repository [^4] https://github.com/kinit-sk/PEFT-Bench. In the future, we plan to use this benchmark to evaluate factors that affect the performance of multi-task PEFT methods.
@@ Acknowledgments
This work was partially funded by the European Union NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the projects No. 09I01-03-V04-00006 and 09I01-03-V04-00068; and by the European Union, under the project LorAI - Low Resource Artificial Intelligence, GA No. 101136646.
@@
We also run PEFT-Bench on a single model – LLaMa3-8B-Instruct. This may limit the generalizability of PEFT-Bench, but, as with language, PEFT methods are not really dependent on the model architecture, and their performance would generally correlate with the model’s performance.
Finally, we evaluate 7 different PEFT methods, with the majority being from the additive PEFT and soft prompt-based category. To keep experiments computationally feasible, we rely on the evaluated PEFT methods with limited hyperparameter tuning. Thanks to PEFT-Factory, the whole benchmark is easily extendable and can add their own PEFT method(s).
@@
The experiments in this paper were conducted with publicly available datasets MNLI, QQP, QNLI, SST2, STS-B, MRPC, RTE, CoLA, MultiRC, ReCoRD, BoolQ, WiC, WSC, CB, COPA, MMLU, PIQA, SIQA, HellaSwag, WinoGrande, OpenBookQA, MathQA, GSM8K, SVAMP, Conala, CodeAlpacaPy, and APPS, citing the original authors. As we were unable to determine the licenses for all used datasets, we have opted to use them in the limited form possible, adhering to the terms of use of the GLUE and SuperGLUE benchmarks. As the datasets are commonly used in other related works and have been published in scientific works that underwent an established review process, we do not check for the presence of any offensive content, as it was already removed by the authors of these publicly available datasets. In addition, we do not utilize any personally identifiable information or offensive content, and we do not perform crowdsourcing in any form for data annotation. To our knowledge, we are not aware of any potential ethical harms or negative societal impacts of our work, apart from the ones related to the field of Machine Learning (i.e., the use of computational resources that are consuming energy and producing heat with indirect CO2 emission production). We follow the license terms for the LLaMa-3-8B-Instruct model we use – all models and datasets allow their use as part of the research. As we transform conditional generation into the classification problem (generating only labels), in most cases, we minimize the problem of generating offensive or biased content.
@@
The experiments in this paper require a significant amount of GPU computing resources as we train and evaluate 1 model over multiple random initializations (5) for different methods (6) and datasets (27). Overall, the experiments, including evaluations (which did not require training, but still used GPU resources for inference) and preliminary experiments (which are not reported in the scope of our work), were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCO2eq/kWh. Approximately 3000 hours of computation were performed on hardware of type A100 PCIe 40GB (TDP of 250W). Total emissions are estimated to be 9.24 kgCO2eq, of which 0 percent were directly offset. Whenever possible, we tried to reduce the computational costs.
@@ Appendix B PSCP Importance Configurations
In this section, we provide an extended example of different configurations for the importance of each cost $\beta$. To evaluate the changes in the magnitude of the importance of each cost $\beta$, we provide results with different magnitudes in Table 9. The order of the methods can change if the performance differences between two methods are similar but differ only in a particular cost.
Based on the results from Table 10, we can see that with the change of importance for each cost, the order of the results can change (e.g., in Table 10(a), the LoRA method achieves the best PSCP score, since we do not penalize the parameters cost as much as the other efficiency factors). Each sub-table represents a different use case, where we emphasize different efficiency factors.
PEFT Method $P_{avg}$ $cost_{p}^{-1}$ $cost_{f}^{-1}$ $cost_{m}^{-1}$ PSCP
IA3 74.7 1.0 0.97 0.77 55.62
Prompt Tuning 50.0 1.0 0.86 0.69 29.49
Prefix Tuning 45.9 0.94 1.0 0.77 33.18
P-Tuning 51.7 0.9 0.86 0.73 29.26
LoRA 80.1 0.97 1.0 0.77 60.08
LNTuning 77.8 1.0 1.0 0.77 60.19
BitFit 75.3 1.0 1.0 0.8 60.1
PEFT Method $P_{avg}$ $cost_{p}^{-2}$ $cost_{f}^{-2}$ $cost_{m}^{-2}$ PSCP
IA3 74.7 1.0 0.94 0.59 41.41
Prompt Tuning 50.0 1.0 0.73 0.47 17.39
Prefix Tuning 45.9 0.88 1.0 0.6 23.98
P-Tuning 51.7 0.82 0.73 0.53 16.56
LoRA 80.1 0.94 1.0 0.6 45.06
LNTuning 77.8 1.0 1.0 0.6 46.57
BitFit 75.3 1.0 1.0 0.64 47.97
PEFT Method $P_{avg}$ $cost_{p}^{-3}$ $cost_{f}^{-3}$ $cost_{m}^{-3}$ PSCP
IA3 74.7 1.0 0.91 0.45 30.83
Prompt Tuning 50.0 1.0 0.63 0.33 10.26
Prefix Tuning 45.9 0.82 1.0 0.46 17.33
P-Tuning 51.7 0.74 0.63 0.39 9.37
LoRA 80.1 0.92 1.0 0.46 33.8
LNTuning 77.8 1.0 1.0 0.46 36.03
BitFit 75.3 1.0 1.0 0.51 38.29
PEFT Method $P_{avg}$ $cost_{p}^{-4}$ $cost_{f}^{-4}$ $cost_{m}^{-4}$ PSCP
IA3 74.7 1.0 0.89 0.35 22.95
Prompt Tuning 50.0 1.0 0.54 0.23 6.05
Prefix Tuning 45.9 0.77 1.0 0.36 12.53
P-Tuning 51.7 0.67 0.54 0.29 5.31
LoRA 80.1 0.89 1.0 0.36 25.35
LNTuning 77.8 1.0 1.0 0.36 27.88
BitFit 75.3 1.0 1.0 0.41 30.56
Table 9: Results with PSCP metric under equal importance of each cost with different magnitudes. Because the performance differences between BitFit and LNTuning are small, the relative importance of each cost changed the ordering of the methods.
@@
BitFit 75.3 1.0 1.0 0.64 47.99
Table 10: Results with PSCP metric with different importance of each cost $\beta$. Notice how the order of the methods changes in Table (a) when we decrease the importance of the cost for the number of trainable parameters $\beta_{p}$.
@@ Appendix C Descriptions of PEFT Methods
BitFit (Bias Fine-tuning) Ben Zaken et al. (2022) fine-tunes only the bias terms of weights in selected modules of transformer architecture. However, in modern LLMs, the bias term is often skipped while designing the model. In our implementation, instead of reusing the pre-trained bias term (as in the original implementation), we replace the linear layer with a new one that also contains the bias term and copy the pre-trained weights.